104g of calcium oxide dissolved in 1 l of water. Calculate the concentration of OH- ions in the solution.
1
Expert's answer
2020-10-26T14:48:13-0400

n(CaO)=m/M=104/56=1.86(mol).

When CaO is dissolved in water, Ca(OH)2 is formed:

CaO+H2O=Ca(OH)2.

It can be seen that number of moles of Ca(OH)2 is equal to the number of moles of CaO.

C(Ca(OH)2)=n/V=1.86/1=1.86(mol/L).

But concentration of OH- ions is two times higher:

C(OH-)=2×1.86=3.72(mol/L).